Calculate the value of the translational partition function of O2 at 300 K for a volume of 1 m3. Submit your answer multiplied by 10−32 (i.e. if you calculate 4.2e33, input 42).
1
Expert's answer
2020-06-02T14:07:39-0400

Solution: mass of O2 is 2*16*1.6606*10-27 kg, where 16 is weight of oxygen atom in a.m.u, 1.6606*10-27 kg is 1 atomic mass unit.

2"\pi"mkBT=2*3.1415*(2*16*1.6606*10-27 kg)*1.3807*10-23 J/K*300 K=1.3829*10-45 J kg.

In the equation written above: kB - Boltzmann constant, J/K.

We find the de Brogile thermal wavelength, "\Lambda", m:

"\Lambda"=h/(2"\pi"mkBT)1/2=6.6262*10-34 J s/(1.3829*10-45)1/2=1.7818*10-11 m.

In the equation written above: h - Planck constant, J s.

We calculate the value of the translation partition function, qtr:

qtr=V/"\Lambda"3=1 m3/(1.7818*10-11 m)3=1.7676*1032. This means that 1.7676*1032 quantum states are thermally accessible to the molecular system.


Answer: 1.8